OpenStudy (anonymous):

the length of a rectangle is five times its width. if the area of the rectangle is 180ft^2, find its perimeter

OpenStudy (anonymous):

i like to draw these out it makes it super easy

OpenStudy (anonymous):

i just need the formula, and then i can do it myself

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

k sec

OpenStudy (anonymous):

180ft^2 = (5w) x (w)

OpenStudy (anonymous):

180ft^2= 5w^2

OpenStudy (anonymous):

36= w^2

OpenStudy (anonymous):

^^)

OpenStudy (anonymous):

btw 6=w and Length= 5(6) P= 2L + 2W

OpenStudy (anonymous):

P= 2(30) + 2(6)